# Bayesian Ego-graph inference for Networked Multi-Agent Reinforcement Learning

### v2

@@ 1 Introduction
Multi-agent reinforcement learning (MARL) has emerged as a powerful framework for sequential decision-making in distributed systems, enabling applications in autonomous driving [1, 2], wireless communication [3, 4], multiplayer games [5, 6], and urban traffic control [7, 8]. A popular training paradigm is centralized training with decentralized execution (CTDE), where a centralized critic leverages global state information to train decentralized  [9, 10, 11].
@@ 4.1 Graph-based Policy with Latent Interaction Structures
4.2 BayesG: Variational Latent Graph Inference for Policy Optimisation
@@
where $\tilde{s}_{i,\tau}$ is the graph-conditioned input for the critic, and $\hat{R}_{i,\tau}^{\pi}$ is the estimated return. The final training objective combines both components across agents:
$\mathcal{L}_{\text{total}}=\sum_{i\in\mathcal{V}}\left(\mathcal{L}_{\mathrm{ELBO}}+\mathcal{L}_{\text{Critic}}\right).$ (11)
This formulation enables decentralized agents to learn both communication structure and policy behavior from local interactions, while adhering to the physical structure of the environment. Full derivations are provided in Appendix A.3.
@@ 6.1 Comparison with Sampling-based Networked MARL
Recent work has explored sampling strategies for scalability in networked MARL, but differs fundamentally in what is sampled, how distributions are determined, and where adaptivity occurs.
Sampling targets and distributions. Qu et al. [25] establish convergence results for networked systems with local dependencies on fixed graphs. Lin et al. [22] sample active links from a predefined distribution $\mathcal{D}$ for $\mu$-decay analysis. Anand and Qu [26] use uniform agent subsampling ($k$ out of $n$) for a global controller with mean-field bounds. In contrast, our method learns context-dependent, per-edge distributions: each agent maintains a learned posterior $q(Z_{i};\phi_{i})$ over its 1-hop edges, optimized jointly with the policy via an ELBO.
Locus of adaptivity. Lin et al. [22] achieve dynamics-level stochasticity through time-varying link sets with multi-hop influence across time. Anand and Qu [26] vary agent subsets for a global controller. Our approach operates at the policy/representation level: each agent samples a task-adaptive 1-hop subgraph for local decision-making.
First-hop inclusion probabilities. Lin et al. [22] use the marginal of $\mathcal{D}$ (predefined, not state-adaptive). Anand and Qu [26] use uniform inclusion ($k/n$) with no edge-specific notion. Our method learns edge-specific probabilities $\Pr[z_{ij}=1]=\sigma(\phi_{ij}(\cdot))$ that are state and trajectory-dependent.
These approaches provide valuable convergence guarantees under structural assumptions. Our contribution is complementary: learning which neighbors to attend to, optimized end-to-end with the policy. Extending their theoretical tools to learned edge masks is a promising future direction. A detailed side-by-side comparison is provided in Table 2 (Appendix H).
@@ Appendix I Limitations
  • Fixed physical topology. Our approach assumes a predefined, static environment graph that restricts the set of possible neighbor interactions. While this models many real-world systems such as traffic networks, it may not generalize to domains with dynamic or learned topology [42].
  • Local observability. Each agent infers its latent communication graph based solely on local observations within its ego-graph. This limits the model’s ability to reason over long-range dependencies that may require more global context [43].
  • Hyperparameter sensitivity. The performance of the learned mask depends on the Gumbel-softmax temperature and sparsity prior. Improper tuning may lead to either under- or over-pruning of edges, potentially degrading coordination [44].
@@
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.

### v3

@@ 1 Introduction
Multi-agent reinforcement learning (MARL) has emerged as a powerful framework for sequential decision-making in distributed systems, enabling applications in autonomous driving [1, 2], wireless communication [3, 4], multiplayer games [5, 6], and urban traffic control [7, 8]. A popular training paradigm is centralized training with decentralized execution (CTDE), where a centralized critic leverages global state information to train decentralized  [9, 10, 11, 12].
@@ 4.1 Graph-based Policy with Latent Interaction Structures
4.2 BayesG: Variational Latent Ego-graph Inference for Policy Optimisation
@@
where $\tilde{s}_{i,\tau}$ is the graph-conditioned input for the critic, and $\hat{R}_{i,\tau}^{\pi}$ is the estimated return. The final training objective combines both components across agents: $\mathcal{L}_{\text{total}}=\sum_{i\in\mathcal{V}}\left(-\mathcal{L}_{\mathrm{ELBO}}+\mathcal{L}_{\text{Critic}}\right).$ This formulation enables decentralized agents to learn both communication structure and policy behavior from local interactions, while adhering to the physical structure of the environment. The ego-graph terminology reflects that each agent independently infers its own local communication neighborhood. Full derivations are provided in Appendix A.3.
@@ 6.1 Comparison with Sampling-based Networked MARL
Recent work has explored sampling strategies for scalability in networked MARL, but differs fundamentally in what is sampled, how distributions are determined, and where adaptivity occurs. (1)Sampling targets and distributions.Qu et al. [26] establish convergence results for networked systems with local dependencies on fixed graphs. Lin et al. [23] sample active links from a predefined distribution $\mathcal{D}$ for $\mu$-decay analysis. Anand and Qu [27] use uniform agent subsampling ($k$ out of $n$) for a global controller with mean-field bounds. In contrast, our method learns context-dependent, per-edge distributions: each agent maintains a learned posterior $q(Z_{i};\phi_{i})$ over its 1-hop edges, optimized jointly with the policy via an ELBO. (2) Locus of adaptivity. Lin et al. [23] achieve dynamics-level stochasticity through time-varying link sets with multi-hop influence across time. Anand and Qu [27] vary agent subsets for a global controller. Our approach operates at the policy/representation level: each agent samples a task-adaptive 1-hop subgraph for local decision-making. (3) First-hop inclusion probabilities. Lin et al. [23] use the marginal of $\mathcal{D}$ (predefined, not state-adaptive). Anand and Qu [27] use uniform inclusion ($k/n$) with no edge-specific notion. Our method learns edge-specific probabilities $\Pr[z_{ij}=1]=\sigma(\phi_{ij}(\cdot))$ that are state and trajectory-dependent. These approaches provide valuable convergence guarantees under structural assumptions. Our contribution is complementary: learning which neighbors to attend to, optimized end-to-end with the policy. Extending their theoretical tools to learned edge masks is a promising future direction. A detailed side-by-side comparison is provided in Table 2 (Appendix H).
@@ Appendix I Limitations
  • Fixed physical topology. Our approach assumes a predefined, static environment graph that restricts the set of possible neighbor interactions. While this models many real-world systems such as traffic networks, it may not generalize to domains with dynamic or learned topology [43, 44, 45].
  • Local observability. Each agent infers its latent communication graph based solely on local observations within its ego-graph. This limits the model’s ability to reason over long-range dependencies that may require more global context [46, 47].
  • Hyperparameter sensitivity. The performance of the learned mask depends on the Gumbel-softmax temperature and sparsity prior. Improper tuning may lead to either under- or over-pruning of edges, potentially degrading coordination [48, 49].